OpenStudy (anonymous):

Never True? Sometimes True? Always True? (x^a)(x^b)=x^ab sqrt x^2=x sqrt x^3=sqrt x^3 x^a/x^b= x^a-b x^0=0 x^2+x^2=x^4

OpenStudy (anonymous):

@jim_thompson5910 can you help with this??

jimthompson5910 (jim_thompson5910):

how far did you get

OpenStudy (anonymous):

I thin i only know one, thats why i really need help

jimthompson5910 (jim_thompson5910):

which one are you referring to

OpenStudy (anonymous):

x^0=0 x^2+x^2=x^4 these TWO i think the first is always true, and the senond is sometimes

jimthompson5910 (jim_thompson5910):

both are incorrect

jimthompson5910 (jim_thompson5910):

here are some rules that you should know by heart http://www.mathsisfun.com/algebra/exponent-laws.html

OpenStudy (anonymous):

yes! thanks , ill take a look at the link

jimthompson5910 (jim_thompson5910):

you're welcome